OpenStudy (dragon123mama):

Will give medal. 1. Solve formula for indicated variable. s=6r^3t , for t 2. the function F(x)=x^2 . translate 3 units to the right and 3 units down. what is the new function? 3. solve the system by substitution 2.25x -3y = -13 3.25x - y = -14

jimthompson5910 (jim_thompson5910):

Hint for #1: If x*y = z, then y = z/x. This happens after you divide both sides by x. The division happens to undo the multiplication. For example, 2*x = 10 will turn into x = 10/2 = 5 after you divide both sides by 2 (to undo the multiplication of 2)

OpenStudy (dragon123mama):

??

jimthompson5910 (jim_thompson5910):

With \[\Large s=6r^3t\] you have \(\Large 6r^3\) times \(\Large t\) on the right side. Agreed?

OpenStudy (dragon123mama):

yeah

jimthompson5910 (jim_thompson5910):

so to isolate t, you undo that multiplication

jimthompson5910 (jim_thompson5910):

divide both sides by 6r^3 to undo the multiplication of 6r^3

jimthompson5910 (jim_thompson5910):

Divide both sides by 6r^3 \[\Large s=6r^3t\] \[\Large \frac{s}{6r^3}=\frac{6r^3t}{6r^3}\] \[\Large \frac{s}{6r^3}=\frac{6r^3}{6r^3}*t\] \[\Large \frac{s}{6r^3}=\frac{\cancel{6r^3}}{\cancel{6r^3}}*t\] \[\Large \frac{s}{6r^3}=t\] \[\Large t=\frac{s}{6r^3}\]

OpenStudy (dragon123mama):

oh i see.. okay thanks.

jimthompson5910 (jim_thompson5910):

The reason why the cancellation occurs is because anything divided by itself is 1. Well anything but 0.

OpenStudy (dragon123mama):

that i know.

OpenStudy (dragon123mama):

do you know any of the other questions?

jimthompson5910 (jim_thompson5910):

Show me how far you got with #2

OpenStudy (dragon123mama):

i didn't get anywhere. i skipped it because i couldn't remember how thoes changes affect the equation.

jimthompson5910 (jim_thompson5910):

I'll give a similar example Let's say we want to shift f(x) = x^3 five units down and two units to the left to go 2 units to the left, we replace every x with x+2 to get f(x) = (x+2)^3 to shift 5 units down, we just subtract 5 at the end, so f(x) = (x+2)^3 - 5

OpenStudy (dragon123mama):

so if we move it 3 units to the right it would be F(x)=(x-3)^2 ?

OpenStudy (dragon123mama):

then 3 units down would be f(x)= (x-3)^2 -3 ?

jimthompson5910 (jim_thompson5910):

`so if we move it 3 units to the right it would be F(x)=(x-3)^2 ?` correct

jimthompson5910 (jim_thompson5910):

`then 3 units down would be f(x)= (x-3)^2 -3 ?` also correct

OpenStudy (dragon123mama):

okay. do you know the last question?

jimthompson5910 (jim_thompson5910):

are you able to solve \(\Large 3.25x - y = -14\) for y? are you able to get y all by itself?

OpenStudy (dragon123mama):

umm hang on.

OpenStudy (dragon123mama):

that would be y=14+3.25x right?

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

now move onto the first equation \(\Large 2.25x -3y = -13\)

OpenStudy (dragon123mama):

i remember i had solved this but she said my asnwer was wrong. so. i must have missed something.

jimthompson5910 (jim_thompson5910):

replace EVERY copy of 'y' with 14+3.25x this works because y is equivalent to this expression (aka they're the same) this is what substitution is all about. Think about your substitute teacher. The substitute teacher is a temporary replacement for your actual teacher. Ideally they would be identical but of course in the real world, they are not

jimthompson5910 (jim_thompson5910):

anyways, we would go from this \[\Large 2.25x -3y = -13\] to this \[\Large 2.25x -3(14+3.25x) = -13\] take note how the y term is now gone. So from here you just solve for x

OpenStudy (dragon123mama):

okay.

OpenStudy (dragon123mama):

2.25x -42 + 9.75x = -13

OpenStudy (dragon123mama):

12x - 42 = -13

OpenStudy (dragon123mama):

am i doing this right?

jimthompson5910 (jim_thompson5910):

you made a mistake when you did 2.25x -42 + 9.75x = -13

jimthompson5910 (jim_thompson5910):

the negative distributes through

OpenStudy (dragon123mama):

oh so 2.25X -42 -9.25X = -13

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (dragon123mama):

then. -7.5x - 42 =-13

OpenStudy (dragon123mama):

-7.5x =29

OpenStudy (dragon123mama):

x==3.8?

OpenStudy (dragon123mama):

-3.8

OpenStudy (dragon123mama):

???

jimthompson5910 (jim_thompson5910):

I'm getting -3.866666667 for x

OpenStudy (dragon123mama):

yeah that is what i got. i siplified.

OpenStudy (dragon123mama):

-3.9 if you round.

jimthompson5910 (jim_thompson5910):

if you're going to round to the nearest tenth, then it would be -3.9

jimthompson5910 (jim_thompson5910):

yes correct